# BSTcontrol

### v2

@@ 5.1.2 Quantum Computing
While moving toward the development of 6G and the era of ubiquitous connectivity, the demand for computational resources is increasing exponentially. To counter this surge of demand, quantum computing has been proposed as a promising solution, offering unprecedented processing capabilities through the principles of quantum mechanics. Quantum computing fundamentally changes how we handle complex computations by utilizing qubits, which can exist in multiple states simultaneously [252]. This unique characteristic enables quantum computers to perform parallel processing on a scale unattainable by classical computers. In the context of WLAM, quantum computing holds the promise of significantly enhancing computational efficiency, optimizing resource allocation, and enabling real-time data processing [253, 232]. In the following we explore various types, applications, and future prospects of quantum computing in wireless communications.
@@
Quantum computing can be categorized into several types, each with unique applications. Blind quantum computing, also known as secure quantum computing with privacy preservation, allows a client to delegate computation tasks to remote quantum computers while keeping the source data confidential [254]. This is achieved by sending transformed qubits to the server, which performs computations and returns temporary results that the client can then convert into the final results, ensuring data privacy throughout the process. Distributed quantum computing involves distributing computational tasks across multiple quantum computers, enhancing data processing speed and decreasing latency. This method is significantly useful for connecting noisy intermediate-scale quantum computers to collaboratively execute complex tasks [233]. Additionally, quantum machine learning (QML) leverages quantum computing for machine learning tasks, enabling novel computing services and applications in the context of 6G networks [234]. Quantum computing offers innovative solutions for meeting the huge computational demands of future wireless communications.
@@
QML, particularly quantum neural networks (QNNs), is emerging as a powerful force for WLAM in 6G. Addressing the escalating computational demands of WLAM, QML offers pathways to enhance efficiency in critical tasks. For instance, QNN-based frameworks, as explored in [235], demonstrate the capability to achieve comparable performance in wireless resource allocation tasks, such as user grouping in NOMA, but with significantly reduced computational complexity compared to classical neural networks. This efficiency gain is crucial for the real-time operation of WLAM in dynamic 6G environments. Furthermore, the integration of QNNs with FL, as studied in [236], leverages quantum teleportation to streamline model aggregation in distributed WLAM deployments, potentially accelerating training and enhancing the scalability of intelligent 6G networks.
@@
The integration of quantum computing with WLAM promises significant advancements in 6G and beyond [237]. As quantum computing progresses, its combination with AI models can vastly improve efficiency and capability in managing the immense data generated by communication networks. Hybrid quantum-classical algorithms could solve complex optimization problems, such as dynamic spectrum management and real-time interference mitigation. Additionally, QNNs combined with advanced AI techniques could enable adaptive, self-optimizing networks that learn and evolve in real time. Quantum-secure communication protocols will enhance data integrity and privacy, meeting the growing computational and connectivity demands of the future. This fusion is poised to revolutionize wireless communication systems, promoting an era of intelligent, efficient, and secure networks.
@@ 5.2.1 Physics-Informed Neural Networks
Figure 16:   Liquid neuron and the ODE modelling, the principle of LTC, CfC and NCP.
@@
While PINNs are often motivated by scientific research challenges, their fundamental strength lies in solving differential equations. Therefore, ordinary differential equation (ODE) a well-established and popular application area for PINNs. ODE describes the fundamental changes in systems over time, linking physical processes with temporal evolution. By modeling how variables change continuously with respect to time, ODE provides a powerful framework for understanding dynamic systems [255]. In the context of neural networks, ODE-based methods leverage this continuous modeling approach to enhance the robustness and efficiency of deep learning models. These networks incorporate ODE to represent the continuous transformation of data, offering a more nuanced and flexible alternative to traditional discrete architectures. The key innovation in ODE-based neural networks is the use of ODE solvers to manage the evolution of hidden states throughout the training and inference processes. This approach enables the network to learn complex, continuous dynamics, which is particularly beneficial for tasks involving temporal or sequential data [256].
@@
Recently, a novel type of ODE-based neural network, known as liquid neural networks (LNNs), has been developed from first principles to address the above shortcomings [257]. Unlike traditional AI models, LNNs are grounded in first principles, which involve deriving properties and behaviors from fundamental natural laws to ensure that the design is both innovative and foundational. Inspired by the adaptive and dynamic nature of biological neural systems, LNNs emulate information transmission mechanisms observed in the nematode Caenorhabditis elegans. As illustrated in Fig. 16, LNNs utilize a nonlinear conductance-based synapse model where stimulation flows from a presynaptic neuron to a postsynaptic neuron. This interaction is described by a first-order ODE, forming the basis of the liquid time-constant neural network (LTC) [258]. By approximating the LTC solution with a closed-form expression that uses a few parameters, the closed-form continuous-time neural network (CfC) [259] is derived, which circumvents the high overhead of traditional ODE solvers. Multiple LTC or CfC can be stacked to form neural circuit policie (NCP), which has stronger expressive power [260]. This approach enables LNNs to emulate the flexibility and resilience of natural neural networks. Unlike static architectures, LNNs can continuously adapt and reorganize in response to new inputs, maintaining high performance and robustness in dynamic and unpredictable environments. This adaptability makes LNNs particularly well-suited for real-world applications where conditions constantly change, such as beam management in dynamic and noisy wireless environments[241, 242]. Furthermore, LNNs are able to decompose complex neural dynamics into interpretable and manageable behavioral patterns. By utilizing techniques such as decision trees to analyze neural strategies, LNNs not only provide clear and logical explanations for decision-making processes but also enhance resilience of the system to disturbances, thereby further improving overall robustness [261].
@@
PINNs are increasingly adopted in intelligent wireless communications to enhance system performance. For instance, PINNs have been instrumental in developing intelligent beamforming schemes. In [238], researchers introduced a PINN based on the discrete dipole approximation method, for code-to-pattern mapping. Complementarily, a deep neural network trained under PINN guidance is used for pattern-to-code mapping. This integration results in a joint scheduling framework that effectively combines horizontal task distribution and vertical computational enhancement to improve overall system performance. Beyond beamforming, PINNs have also demonstrated successful application in channel prediction. The SCGnet scheme, presented in [239], exemplifies this by innovatively modeling channel prediction as an ODE problem grounded in electromagnetic wave propagation physics and inspired by Neural ODEs. This data-efficient approach requires only historical data for training and minimal fresh measurements for prediction, showcasing superior performance in mobile channel representation, learning, and prediction. Furthermore, radiomap estimation benefits from PINN methodologies, as demonstrated by the physics-informed machine learning methods proposed in [240] for high-resolution radiomap reconstruction from sparse samples. Their findings underscore the promising synergy between data-driven AI and model-based radio propagation understanding for this task.
@@
Secondly, PINNs can bring enhanced explainability. Traditional DNNs, while powerful, are often criticized for their ”black-box” nature, making it difficult to understand their decision-making processes. This lack of explainability can be a barrier to trust and deployment, especially in critical communication infrastructure. PINNs, by incorporating known physical laws as constraints or guiding principles within their architecture, inherently offer enhanced explainability. The learned network is not solely driven by data patterns but also by the imposed physical relationships. This allows researchers and engineers to gain insights into why a PINN is making specific predictions or decisions. By examining the learned parameters and how they interact with the embedded physical models, we can better understand the underlying wireless phenomena being captured by the network. This improved interpretability facilitates debugging, validation, and ultimately, greater confidence in the reliability and robustness of PINN-driven wireless communication systems.
@@
Looking ahead, the integration of PINNs with the burgeoning field of large AI models holds immense promise for revolutionizing wireless communication. As wireless systems become increasingly complex, demanding ultra-high performance and adaptability in dynamic environments, the limitations of purely data-driven large models in terms of generalization, explainability, and sample efficiency become more pronounced. PINNs offer a compelling solution by injecting fundamental physical principles into these large models, potentially leading to a new generation of wireless AI that is not only powerful but also interpretable and robust. It is anticipated that future 6G networks leveraging PINN-enhanced large models for tasks like massive multi-input multi-output (MIMO) beamforming, intelligent spectrum management, and network optimization, would achieve unprecedented levels of efficiency and reliability. By grounding large models in the known physics of wireless propagation and system behavior, we can construct more trustworthy, resource-efficient, and ultimately, more capable wireless networks that can seamlessly navigate the complexities of future communication landscapes.
@@ 5.2.2 Hyper-networks
To address the diverse needs of various applications and users, Hyper-networks (HNs) represent a novel neural network architecture designed to dynamically generate the weights of another network [262]. Traditional neural networks require re-training or fine-tuning to accommodate specific user needs, which is both expensive and time-consuming. HNs overcome this limitation by producing whole or partial weights based on input conditions, thus enabling adaptive and flexible model behavior without the need for re-training or fine-tuning. This dynamic generation process allows for rapid adjustments, making HNs highly suitable for complex and versatile environments. In the following we delve into the applications, benefits, and recent developments of HNs in large AI models for wireless communication.
@@
HNs provide substantial benefits for large AI models in wireless communication through dynamic resource allocation, as shown in Fig. 17. The HNs-based resource allocation framework comprises two parts: the resource allocation component and the adaptive control component. The large AI model within the resource allocation component manages resource allocation in wireless systems based on inputs such as CSI. Meanwhile, they can customize the full or partial weights of the large AI model using the input conditional vector, such as user weights and location. Specifically, they can generate adaptive beamforming vectors and configurations based on user weights in RIS-assisted systems, thus enhancing efficiency without retraining [243]. Additionally, they improve non-stationary channel prediction by continuously updating neural network parameters to adapt to changing conditions, which increases CSI prediction accuracy and spectral efficiency [263]. Combined with RNNs, HNs leverage uplink-downlink reciprocity to enhance channel estimation and beamforming performance in frequency division duplex massive MIMO systems [264]. This synergy facilitates high-performance processing in environments with limited hardware capabilities, optimizing resource allocation and operational efficiency in wireless applications. HNs reduce the need for extensive retraining by dynamically generating network weights, making them particularly well-suited for rapidly changing environments and diverse user requirements. Their integration with large AI models supports scalable and efficient solutions for modern wireless networks.
@@
Recent developments in the application of HNs in large AI models showcase innovative approaches to enhancing model efficiency and performance. [265] introduces a novel approach that uses HNs to transform task instructions into parameter-efficient modules, significantly reducing computational costs while improving performance by up to 25% compared to state-of-the-art methods. Concurrently, [244] offers a new framework for generating large AI models tailored to specific tasks or data descriptions, achieving up to 270x speed improvements over traditional fine-tuning approaches. These developments not only highlight the potential for more efficient and effective model inference but also pave the way for future innovations in wireless communication. The integration of HNs and user-customized models, as exemplified above, promises to enhance the adaptability, scalability, and performance of large AI models. By leveraging these advancements, future wireless AI systems can achieve more tailored solutions for diverse communication needs, ultimately advancing the field towards more intelligent and resource-efficient technologies.
@@ 5.2.3 Next-Gen Sequence Modeling Networks
The emergence of 6G communication systems demands advanced AI models capable of processing large-scale sequential data with high efficiency. While Transformer-based architectures have demonstrated high performance, their quadratic complexity with respect to sequence length poses scalability challenges. To address this, next-generation sequence modeling networks, including Mamba [245], receptance weighted key value (RWKV) [246], and test-time training (TTT) [247], have been developed, offering improved computational efficiency and adaptability. This subsection examines these architectures, detailing their mechanisms, advantages, and potential applications in large AI model empowered 6G networks.
@@
In 6G systems, Mamba stands out as a powerful solution for reducing the computational overhead of large AI models in communication systems. Its linear complexity allows it to efficiently process vast sequential data, a critical advantage for large models handling extensive datasets without incurring excessive resource costs. This efficiency is particularly valuable for tasks like dynamic channel estimation and adaptive beamforming. Its real-time processing capabilities enable rapid adaptation to changing conditions, which is essential for high-performance, low-latency communication. Besides, Mamba has demonstrated effectiveness in related areas including wireless sensing. A concrete example is the SenseMamba architecture [266], illustrated in Fig. 18, which applies selective SSM principles within specifically designed lightweight blocks to achieve efficient human sensing using wireless signals. This demonstrates its potential not only for core communication tasks like channel estimation but also for enabling sophisticated, low-overhead sensing applications crucial for future intelligent environments. Its real-time processing capabilities and ability to manage complex temporal dependencies make it particularly suitable for these dynamic scenarios. Furthermore, Mamba is especially beneficial for large AI models for its ability to manage infinite contexts without overwhelming resource demands. Large models often require deep, context-rich understanding to perform effectively, and Mamba delivers this capability seamlessly, decreasing the overhead that typically burdens such systems.

### v4

@@ 5.1.2 Quantum Computing
While moving toward the development of 6G and the era of ubiquitous connectivity, the demand for computational resources is increasing exponentially. To counter this surge of demand, quantum computing has been proposed as a promising solution, offering unprecedented processing capabilities through the principles of quantum mechanics. Quantum computing fundamentally changes how we handle complex computations by utilizing qubits, which can exist in multiple states simultaneously. This unique characteristic enables quantum computers to perform parallel processing on a scale unattainable by classical computers. In the context of WLAM, quantum computing holds the promise of significantly enhancing computational efficiency, optimizing resource allocation, and enabling real-time data processing [wang2022quantum, botsinis2018quantum]. In the following we explore various types, applications, and future prospects of quantum computing in wireless communications.
@@
Quantum computing can be categorized into several types, each with unique applications. Blind quantum computing, also known as secure quantum computing with privacy preservation, allows a client to delegate computation tasks to remote quantum computers while keeping the source data confidential [barz2012demonstration]. This is achieved by sending transformed qubits to the server, which performs computations and returns temporary results that the client can then convert into the final results, ensuring data privacy throughout the process. Distributed quantum computing involves distributing computational tasks across multiple quantum computers, enhancing data processing speed and decreasing latency. This method is significantly useful for connecting noisy intermediate-scale quantum computers to collaboratively execute complex tasks [buhrman2003distributed]. Additionally, quantum machine learning (QML) leverages quantum computing for machine learning tasks, enabling novel computing services and applications in the context of 6G networks [narottama2023quantum]. Quantum computing offers innovative solutions for meeting the huge computational demands of future wireless communications.
@@
QML, particularly quantum neural networks (QNNs), is emerging as a powerful force for WLAM in 6G. Addressing the escalating computational demands of WLAM, QML offers pathways to enhance efficiency in critical tasks. For instance, QNN-based frameworks, as explored in [narottama2021quantum], demonstrate the capability to achieve comparable performance in wireless resource allocation tasks, such as user grouping in non-orthogonal multiple access, but with significantly reduced computational complexity compared to classical neural networks. This efficiency gain is crucial for the real-time operation of WLAM in dynamic 6G environments. Furthermore, the integration of QNNs with FL, as studied in [narottama2023federated], leverages quantum teleportation to streamline model aggregation in distributed WLAM deployments, potentially accelerating training and enhancing the scalability of intelligent 6G networks.
@@
The integration of quantum computing with WLAM promises significant advancements in 6G and beyond [chen2021quantum]. As quantum computing progresses, its combination with AI models can vastly improve efficiency and capability in managing the immense data generated by communication networks. Hybrid quantum-classical algorithms could solve complex optimization problems, such as dynamic spectrum management and real-time interference mitigation. Additionally, QNNs combined with advanced AI techniques could enable adaptive, self-optimizing networks that learn and evolve in real time. Quantum-secure communication protocols will enhance data integrity and privacy, meeting the growing computational and connectivity demands of the future. This fusion is poised to revolutionize wireless communication systems, promoting an era of intelligent, efficient, and secure networks.
@@ 5.2.1 Physics-Informed Neural Networks
Figure 17:   Liquid neuron and the ODE modeling, the principle of LTC, CfC and NCP.
@@
While PINNs are often motivated by scientific research challenges, their fundamental strength lies in solving differential equations. Therefore, ordinary differential equation (ODE) a well-established and popular application area for PINNs. ODE describes the fundamental changes in systems over time, linking physical processes with temporal evolution. By modeling how variables change continuously with respect to time, ODE provides a powerful framework for understanding dynamic systems [ODE]. In the context of neural networks, ODE-based methods leverage this continuous modeling approach to enhance the robustness and efficiency of deep learning models. These networks incorporate ODE to represent the continuous transformation of data, offering a more nuanced and flexible alternative to traditional discrete architectures. The key innovation in ODE-based neural networks is the use of ODE solvers to manage the evolution of hidden states throughout the training and inference processes. This approach enables the network to learn complex, continuous dynamics, which is particularly beneficial for tasks involving temporal or sequential data [lstm-ode].
@@
Recently, a novel type of ODE-based neural network, known as liquid neural networks (LNNs), has been developed from first principles to address the above shortcomings [ltc]. Unlike traditional AI models, LNNs are grounded in first principles, which involve deriving properties and behaviors from fundamental natural laws to ensure that the design is both innovative and foundational. Inspired by the adaptive and dynamic nature of biological neural systems, LNNs emulate information transmission mechanisms observed in the nematode Caenorhabditis elegans. As illustrated in Fig. 17, LNNs utilize a nonlinear conductance-based synapse model where stimulation flows from a presynaptic neuron to a postsynaptic neuron. This interaction is described by a first-order ODE, forming the basis of the liquid time-constant neural network (LTC) [ltc]. By approximating the LTC solution with a closed-form expression that uses a few parameters, the closed-form continuous-time neural network (CfC) [cfc] is derived, which circumvents the high overhead of traditional ODE solvers. Multiple LTC or CfC can be stacked to form neural circuit policie (NCP), which has stronger expressive power [ncps]. This approach enables LNNs to emulate the flexibility and resilience of natural neural networks. Unlike static architectures, LNNs can continuously adapt and reorganize in response to new inputs, maintaining high performance and robustness in dynamic and unpredictable environments. This adaptability makes LNNs particularly well-suited for real-world applications where conditions constantly change, such as beam management in dynamic and noisy wireless environments[ncps_beamforming, zhu2024robustcontinuoustimebeamtracking, zhu2025liquidneuralnetworksnextgeneration]. Furthermore, LNNs are able to decompose complex neural dynamics into interpretable and manageable behavioral patterns. By utilizing techniques such as decision trees to analyze neural strategies, LNNs not only provide clear and logical explanations for decision-making processes but also enhance resilience of the system to disturbances, thereby further improving overall robustness [Interpretability].
@@
PINNs are increasingly adopted in intelligent wireless communications to enhance system performance. For instance, PINNs have been instrumental in developing intelligent beamforming schemes. In [li2022intelligent], researchers introduced a PINN based on the discrete dipole approximation method, for code-to-pattern mapping. Complementarily, a deep neural network trained under PINN guidance is used for pattern-to-code mapping. This integration results in a joint scheduling framework that effectively combines horizontal task distribution and vertical computational enhancement to improve overall system performance. Beyond beamforming, PINNs have also demonstrated successful application in channel prediction. The SCGnet scheme, presented in [10351043], exemplifies this by innovatively modeling channel prediction as an ODE problem grounded in electromagnetic wave propagation physics and inspired by Neural ODEs. This data-efficient approach requires only historical data for training and minimal fresh measurements for prediction, showcasing superior performance in mobile channel representation, learning, and prediction. Furthermore, radiomap estimation benefits from PINN methodologies, as demonstrated by the physics-informed machine learning methods proposed in [10634040] for high-resolution radiomap reconstruction from sparse samples. Their findings underscore the promising synergy between data-driven AI and model-based radio propagation understanding for this task.
@@
Secondly, PINNs can bring enhanced explainability. Traditional dense neural networks, while powerful, are often criticized for their ”black-box” nature, making it difficult to understand their decision-making processes. This lack of explainability can be a barrier to trust and deployment, especially in critical communication infrastructure. PINNs, by incorporating known physical laws as constraints or guiding principles within their architecture, inherently offer enhanced explainability. The learned network is not solely driven by data patterns but also by the imposed physical relationships. This allows researchers and engineers to gain insights into why a PINN is making specific predictions or decisions. By examining the learned parameters and how they interact with the embedded physical models, we can better understand the underlying wireless phenomena being captured by the network. This improved interpretability facilitates debugging, validation, and ultimately, greater confidence in the reliability and robustness of PINN-driven wireless communication systems.
@@
Integration Mechanism with WLAMs
The technical integration of PINN principles into a WLAM architecture is primarily achieved through the modification of the training objective via a hybrid loss function. A conventional WLAM, such as a large Transformer network for channel prediction, is trained to minimize a data-driven loss, typically the MSE between its predictions and the ground-truth data $\mathcal{L}_{data}$. A PINN-enhanced WLAM introduces a second component to this objective: a physics-residual loss $\mathcal{L}_{phys}$.
This $\mathcal{L}_{phys}$ term is derived from the governing physical laws of the wireless environment, often expressed as partial differential equations (PDEs) like Maxwell’s equations. The output of WLAM (e.g., the predicted CSI) is treated as a function and substituted into the PDE. The residual of this equation forms the physics-based loss, and the total loss function for training the WLAM becomes a weighted sum: $\mathcal{L}_{total}=\lambda\mathcal{L}_{data}+(1-\lambda)\mathcal{L}_{phys}$. By minimizing this combined loss, the WLAM is forced to find a representation that is not only accurate with respect to the training samples but is also consistent with the fundamental principles of wave propagation. This mechanism is the root of its enhanced data efficiency and generalization capabilities, as the physical loss acts as a powerful regularizer, guiding the model towards a valid solution even with sparse data.
@@
Looking ahead, the integration of PINNs with the burgeoning field of large AI models holds immense promise for revolutionizing wireless communication. As wireless systems become increasingly complex, demanding ultra-high performance and adaptability in dynamic environments, the limitations of purely data-driven large models in terms of generalization, explainability, and sample efficiency become more pronounced. PINNs offer a compelling solution by injecting fundamental physical principles into these large models, potentially leading to a new generation of wireless AI that is not only powerful but also interpretable and robust. It is anticipated that future 6G networks leveraging PINN-enhanced large models for tasks like massive MIMO beamforming, intelligent spectrum management, and network optimization, would achieve unprecedented levels of efficiency and reliability. By grounding large models in the known physics of wireless propagation and system behavior, we can construct more trustworthy, resource-efficient, and ultimately, more capable wireless networks that can seamlessly navigate the complexities of future communication landscapes.
@@ 5.2.2 Hyper-networks
To address the diverse needs of various applications and users, Hyper-networks (HNs) represent a novel neural network architecture designed to dynamically generate the weights of another network [ha2017hypernetworks]. Traditional neural networks require re-training or fine-tuning to accommodate specific user needs, which is both expensive and time-consuming. HNs overcome this limitation by producing whole or partial weights based on input conditions, thus enabling adaptive and flexible model behavior without the need for re-training or fine-tuning. This dynamic generation process allows for rapid adjustments, making HNs highly suitable for complex and versatile environments. In the following we delve into the applications, benefits, and recent developments of HNs in large AI models for wireless communication.
@@
HNs provide substantial benefits for large AI models in wireless communication through dynamic resource allocation. For instance, they can generate adaptive beamforming vectors and configurations based on user weights in reconfigurable intelligent surfaces-assisted systems [10238370], thus enhancing efficiency without retraining [abouamer2023flexible]. Additionally, they improve non-stationary channel prediction by continuously updating neural network parameters to adapt to changing conditions, which increases CSI prediction accuracy and spectral efficiency [liu2024hypernetwork]. Combined with RNNs, HNs leverage uplink-downlink reciprocity to enhance channel estimation and beamforming performance in frequency division duplex massive MIMO systems [liu2022learning]. This synergy facilitates high-performance processing in environments with limited hardware capabilities, optimizing resource allocation and operational efficiency in wireless applications. HNs reduce the need for extensive retraining by dynamically generating network weights, making them particularly well-suited for rapidly changing environments and diverse user requirements. Their integration with large AI models supports scalable and efficient solutions for modern wireless networks.
Integration Mechanism with WLAMs
The integration of HNs with WLAMs is achieved through a dynamic, two-level architectural design. In this framework, the massive WLAM serves as the ‘main network’ for executing the primary task (e.g., resource allocation), while a significantly smaller HN acts as a ‘controller network’, as shown in Fig. 18. The key mechanism is that the HN does not process the primary task data (e.g., CSI); instead, it takes contextual metadata (such as user QoS profiles, device identities, or specific task instructions) as input.
The output of the HN is not a prediction, but rather the weights for specific, adaptable parts of the main WLAM. Most effectively, an HN can be trained to generate the low-rank matrices for PEFT modules, such as LoRA adapters within the WLAM. This creates a highly efficient workflow: when the context changes (e.g., a new user with different requirements connects), only the small HN needs to perform a forward pass to generate a new, customized set of adapter weights. These weights are then injected into the frozen WLAM, instantly tailoring its behavior without the need for any backpropagation through the multi-billion parameter model. This mechanism provides a concrete pathway for achieving rapid and low-cost personalization and task adaptation for WLAMs.
@@
Recent developments in the application of HNs in large AI models showcase innovative approaches to enhancing model efficiency and performance. [ivison2022hint] introduces a novel approach that uses HNs to transform task instructions into parameter-efficient modules, significantly reducing computational costs while improving performance by up to 25% compared to state-of-the-art methods. Concurrently, [tang2024modelgpt] offers a new framework for generating large AI models tailored to specific tasks or data descriptions, achieving up to 270x speed improvements over traditional fine-tuning approaches. These developments not only highlight the potential for more efficient and effective model inference but also pave the way for future innovations in wireless communication. The integration of HNs and user-customized models, as exemplified above, promises to enhance the adaptability, scalability, and performance of large AI models. By leveraging these advancements, future wireless AI systems can achieve more tailored solutions for diverse communication needs, ultimately advancing the field towards more intelligent and resource-efficient technologies.
@@ 5.2.3 Next-Gen Sequence Modeling Networks
The emergence of 6G communication systems demands advanced AI models capable of processing large-scale sequential data with high efficiency. While Transformer-based architectures have demonstrated high performance, their quadratic complexity with respect to sequence length poses scalability challenges. To address this, next-generation sequence modeling networks, including Mamba [gu2024mamba], receptance weighted key value (RWKV) [peng-etal-2023-rwkv], and test-time training (TTT) [sun2024learninglearntesttime], have been developed, offering improved computational efficiency and adaptability. This subsection examines these architectures, detailing their mechanisms, advantages, and potential applications in large AI model empowered 6G networks.
@@
In 6G systems, Mamba stands out as a powerful solution for reducing the computational overhead of large AI models in communication systems. Its linear complexity allows it to efficiently process vast sequential data, a critical advantage for large models handling extensive datasets without incurring excessive resource costs. This efficiency is particularly valuable for tasks like dynamic channel estimation and adaptive beamforming. Its real-time processing capabilities enable rapid adaptation to changing conditions, which is essential for high-performance, low-latency communication. Besides, Mamba has demonstrated effectiveness in related areas including wireless sensing. A concrete example is the SenseMamba architecture [sensemamba], illustrated in Fig. 19, which applies selective SSM principles within specifically designed lightweight blocks to achieve efficient human sensing using wireless signals. This demonstrates its potential not only for core communication tasks like channel estimation but also for enabling sophisticated, low-overhead sensing applications crucial for future intelligent environments. Its real-time processing capabilities and ability to manage complex temporal dependencies make it particularly suitable for these dynamic scenarios. Furthermore, Mamba is especially beneficial for large AI models for its ability to manage infinite contexts without overwhelming resource demands. Large models often require deep, context-rich understanding to perform effectively, and Mamba delivers this capability seamlessly, decreasing the overhead that typically burdens such systems.
@@
Integration Mechanism with WLAMs
For next-gen sequence models, the integration with WLAMs is a direct architectural substitution at the core building-block level. Many proposed WLAMs for sequential wireless tasks (e.g., time-series channel prediction, traffic forecasting) are based on the powerful but computationally expensive Transformer architecture. The integration mechanism involves replacing the self-attention blocks of Transformer, which are the source of its $O(n^{2})$ complexity, with the more efficient linear-time blocks from these next-gen models.
For example, in a Mamba-based WLAM, the self-attention layer is swapped with a SSM block. This fundamentally changes how the model processes sequential information. Instead of comparing every element of the sequence with every other element, the SSM maintains a compact hidden state that is updated recurrently, allowing for $O(n)$ complexity. This architectural change enables a WLAM to process significantly longer sequences of wireless data (e.g., capturing channel coherence over extended periods) in real-time. This is not merely an efficiency gain; it allows the WLAM to learn long-range temporal dependencies in the wireless channel that are computationally intractable for a standard Transformer, leading to potentially more accurate and robust predictions.
@@
5.3 Tensions and Synergies
While the aforementioned technologies offer individual advantages, their practical integration into 6G systems involves complex inter-dependencies. The following analysis reveals both inherent tensions and potential synergies that define the future architectural roadmap.
5.3.1 Tensions: Hardware Maturity and Algorithmic Efficiency
A fundamental tension exists between emerging computing paradigms and neural architectures regarding deployment readiness. While computing paradigms like quantum computing and HDC offer theoretical energy efficiency superior to classical AI, they suffer from low hardware maturity and a lack of standardized toolchains. In contrast, emerging neural architectures such as Mamba and TTT are designed to run efficiently on existing GPUs but hit the “memory wall” on resource-limited edge devices. Consequently, a critical trade-off arises between immediate compatibility and hardware maturity. System designers must balance the deployment of algorithmic innovations that run on existing processors against the pursuit of disruptive computing paradigms, which offer theoretical superiority but currently lack the necessary specialized hardware infrastructure.
5.3.2 Synergies: Physics-Informed Efficiency and Dynamic Adaptation
Conversely, significant synergies emerge when combining these technologies. First, the integration of PINNs with next-generation sequence models creates a powerful duality. While Mamba and RWKV excel at processing long-sequence data with linear complexity, they often lack physical interpretability. By embedding PINN-derived physical constraints into the state-space transitions of Mamba, one can construct “Physics-Informed State Space Models.” These hybrid architectures may achieve high-precision channel prediction with significantly fewer training samples by leveraging physical laws to guide the long-context extrapolation. Second, Hyper-networks offer a synergistic solution to large foundation models. Instead of retraining a massive WLAM for every new user scenario, a lightweight Hyper-network can dynamically generate task-specific adapters for the frozen foundation model. This combination is promising to resolve the conflict between the need for generalist knowledge and the requirement for rapid, user-specific personalization in dynamic wireless environments.